OpenStudy (anonymous):

a force of 2.4N is exerted on a -1.8UC charge in downward direction. what is the magnitude and direction of the electric field at this point?

OpenStudy (cj7529):

[guessing the "-1.8UC" is actually -1.8 micro coulomb] Using the equation F=Eq where F = force E= Electric field strength q= charge E = F/q E= (2.4)/(-1.8x10^-6) E=1.33x10^6 N/C or T Also, if the force is exerted downwards, the field points upwards. This is because the direction of an electric field is defined to be the direction of the force felt by a POSITIVE charge placed within it